OpenStudy (anonymous):

Find the derivative of f(x) = -6/x at x = 12.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

first set up the expression (f(x+h) - f(x)) / h

OpenStudy (anonymous):

f'(x)=6/x^2 f'(12)=6/(12^2)

OpenStudy (anonymous):

\[f(x)=-\frac{ 6 }{ x }\] at x=12

OpenStudy (anonymous):

How would I change it to that?

OpenStudy (phi):

first , the x=12 is used at the last step. second, f(x+h) means replace x in your function definition with x+h

OpenStudy (phi):

fill in what you can in: \[ \lim_{h \rightarrow 0} \frac{ f(x+h) - f(x)}{h} \]

OpenStudy (anonymous):

I understood you to ask for the derivative of f(x) = -6/x = -6x^(-1). There's a differentiation rule for powers of x that gives (d/dx)x^(-1) = -1x(^-2). This is what I used. (-6)(-1x^(-2)) = 6/x^2 Or, of course you can take phi's approach which is lower-level and always works.

OpenStudy (phi):

@BillBell yes, @This_Is_Batman is using the fundamental def of the derivative for his problems.

OpenStudy (phi):

for this problem, it might be less confusing to write it as \[ \lim_{h \rightarrow 0} \frac{1}{h} ( f(x+h) - f(x) ) \] It's the same thing, but less confusing (maybe?)

OpenStudy (phi):

so far you have \[ f(x)=-\frac{ 6 }{ x } \] you also need f(x+h)

OpenStudy (anonymous):

Would I do -6/x + -6/h?

OpenStudy (anonymous):

Sorry @phi , I was starving to death..

OpenStudy (phi):

are you asking what is f(x+h) ? write down f(x) everywhere you see x, erase it and put in (x+h)

OpenStudy (anonymous):

That's what I mean. I have -6/x, so wouldn't I do -6/x+h, which is the same as -6/x -6/h?

OpenStudy (phi):

you can't split -6/(x+h) (you might be thinking of (x+h)/6 = x/6 + h/6 )

OpenStudy (anonymous):

Oh, my bad. :p

OpenStudy (anonymous):

@phi: ok, understood

OpenStudy (anonymous):

So I just have to figure out -(6/(x+h))-(-(6/x))?

OpenStudy (phi):

try to be complete \[ \lim_{h \rightarrow 0} \frac{1}{h} \left( \frac{-6}{x+h} - \frac{-6}{x} \right) \] to add/subtract fractions you need a common denominator.

OpenStudy (anonymous):

So now it's: \[\lim_{h \rightarrow 0}(\frac{ -12h }{ x+h })\]

OpenStudy (phi):

how did you get that ? I would multiply the first fraction by x/x and the second fraction by (x+h)/(x+h) can you do that ?

OpenStudy (anonymous):

Wow, I am retarded. I see what we are doing. It should be: \[\frac{ -12x - 6h }{ x^2 +h}\]

OpenStudy (phi):

x+h and x are not the same number (so not the common denominator) we don't have a way to "add h" to the second x to make it x+h (while keeping the fraction unchanged) all we know how to do is multiply: x(x+h) is the common denominator

OpenStudy (anonymous):

I meant xh, not just h on the denominator.

OpenStudy (anonymous):

The derivative of f is the function whose value at x is the limit

OpenStudy (phi):

closing in on it. but you have a - (-6) so things cancel also x(x+h) is not x^2 + h

OpenStudy (anonymous):

Oh, so just -6h/(x^2+xh)...

OpenStudy (anonymous):

Which would simplify down to -6/(x^2+x), right?

OpenStudy (phi):

double check your algebra. I get +6h up top

OpenStudy (anonymous):

Yup, my bad. 6/(x^2+x)

OpenStudy (anonymous):

I have to get rid of the x...

OpenStudy (phi):

try be be more careful \[ \lim_{h \rightarrow 0} \frac{1}{h} \left( \frac{-6}{x+h} - \frac{-6}{x} \right) \\ \lim_{h \rightarrow 0} \frac{1}{h} \left( \frac{6h}{x^2+xh)} \right) \]

OpenStudy (phi):

notice that 1/h times h cancel and you get \[ \lim_{h \rightarrow 0} \left( \frac{6}{x^2+xh} \right) \] now "take the limit"

OpenStudy (anonymous):

I thought the h on the bottom would go away?

OpenStudy (phi):

Do you see what I posted ? after you put the fractions over a common denominator you get \[ \lim_{h \rightarrow 0} \frac{1}{h} \left( \frac{6h}{x^2+xh)} \right) \] at this point we cannot (yet) let h=0 because h/h would be undefined the next step is simplify h/h to get \[ \lim_{h \rightarrow 0} \left( \frac{6}{x^2+xh)} \right) \] now let h approach 0. the xh "goes away"

OpenStudy (phi):

and the answer is 6/x^2

OpenStudy (anonymous):

Oh, so it goes down to 6/x^2. Which would be 1/24.